slope midpoint distance for (-5,6) AND (1, -8)
1
Expert's answer
2012-10-25T09:41:32-0400
slope=(y2-y1)/(x2-x1)=(6-(-8))/(-5-1)=-14/6=-7/3
midpoint
x=(x1+x2)/2=(-5+1)/2=-2
y=(y1+y2)/2=(6+(-8))/2=-1
distance=sqrt(
(x1-x2)^2+(y1-y2)^2)=sqrt(36+14^2)=sqrt(232)=2sqrt(58)